A sample of hydrogen gas (H2) at 1.00 atm and 273 K has a volume of 225 mL. How many moles of hydrogen gas are there? 


1
Expert's answer
2021-02-17T07:48:45-0500

As we know that PV= nRT for an ideal gas. Now n=PV/RT

P = 1.00 atm, V= 0.225 L, R = 0.082 L. atm. K-1. mol-1 , T= 273 K

therefore n=(1x 0.225)/(0.082x 273) mol =0.01005 mol

So there are 0.01005 moles of hydrogen gas.